How many liters of 4 M solution can be made using 88 moles of LiBr?Given:Work:M = _______mol = _______Liters = ______
Answer #1

\(M=\frac{n}{V}\)

\(M=\) molarity \(=4.0\)

\(V=\) volume of solution in liters \(=?\)

\(n=\) number of moles of \(L i B r=88\)

\(V=\frac{n}{M}=\frac{88}{4}\)

\(V=22\) liters

answered by: Gavin